# Photodissociation spectra of single trapped CaOH molecular ions

### v1

@@ Acknowledgements
This research was funded by ERC-2020-STG 948893, ESQ Discovery project SDEF: State-dependent force spectroscopy for trapped ions, and FWF 1000 Ideas project TAI-798. The authors would like to acknowledge AQT for support with the ion trap and vacuum pumps, in particular Daniel Nigg and Georg Jacob; Markus Teller for assistance with vacuum work; electronics workshop technicians Kilian Prokop and Wolfgang Kuen for lab infrastructure support; Milan Oncak for discussions on quantum chemistry calculations; Christian Marciniak for general discussions; Marco Valentini for Ca lasers implementation; and the Quantum Optics & Spectroscopy Group and associated ion trapping groups at the Universität Innsbruck for general assistance.
Author Contributions Z.W., S.W., B.F., and P.S. designed the experiment. Z.W., S.W., and B.F. carried out the dissociation experiment and analyzed the data. M.I.M., E.M., and B.F. performed the characterization of the OPA system. Z.W., S.W., and B.F. contributed to the manuscript. All authors contributed to the experimental setup and reviewed the manuscript. B.F. and P.S. supervised the project.
@@
The data and analysis code that support the findings of this study are openly available at https://github.com/quantum-molecules/dissociation_analysis.
@@ A.3 Memory of the photodissociation process
We can therefore also estimate the mean dissociation time using the estimated dissociation probability for fixed times. These probabilities are then fitted to the exponential distribution CDF as indicated in Fig. 7, e.g., for nm and nm. We estimate the photodissociation cross sections via this analysis for all data and compare them to the cumulative dissociation time, as shown in Fig. 8. We find that the cross sections obtained from both methods agree within the uncertainties which indicates that the single- and two-photon dissociation processes are memoryless.
@@ Appendix B Experimental details
B.1 Ca Lasers and trap parameters
@@ Appendix C Ion crystal configuration identification
The photodissociation experiment requires real-time evaluation of the ion crystal configuration such that the number of fluorescing (bright) Ca and non-fluorescing (dark) molecular ions can be monitored. Knowledge of the ion crystal configuration allows a script to run automatic measurements and take actions such as (re)loading the trap or emptying it if more than one dark ion is created. The ion crystal configuration is identified using only image data of the ions taken by an imaging system with magnification using an Andor iXon+ 885 camera with a pixel size of 8 µm. For this process, a region of interest (ROI) in the image centered around the ions is converted into two normalized data sets: the column sum, where the ROI is summed along the columns of the image array; and the row sum, where the summation is done along the rows (see Fig. 11). All analysis to identify the ion crystal configuration is based on these two data sets.
@@
The number of bright ions is counted from the row sum via a peak finder algorithm, which can determine the number and positions of bright ions without requiring a fluorescence-based threshold. This algorithm identifies the largest amplitude difference in a set of sorted local maxima (peaks) of row sum and uses this to classify peaks into two distinct groups: large-amplitude peaks originating from bright ions and small-amplitude peaks due to noise. Knowledge of the positions of the bright ion peaks can be utilized to identify the presence of dark ions for, e.g., determining the trap center position, as it requires an ion crystal of solely bright ions.
In order to count dark ions, a template fitting procedure was performed to identify the ion crystal configuration. In this process, template sets for different ion numbers and configurations are compared to the column sum data set. Template sets are generated by the following scheme. The desired ion configuration in a template for ions is given by the vector . The value of is assigned to for a bright Ca ion and for a dark molecular ion, for , e.g., the ion crystal configuration template which matches the row sum shown in Fig. 11 is . Then the sum of Gaussian peaks
@@
Here denotes the numerically calculated positionsJames (1998) of the ions with respect to the calibrated trap center position , is the waist of the Gaussian peaks (taken from the fit presented in Fig. 11), and is an offset of the trap center position. The start and end positions, and , of the -th binning interval are chosen such that the binning and length of the template match those of the measured column sum data set. is the binned template of ions for the configuration . To find the particular ion crystal configuration, a large set of templates is created for all reasonable ion numbers and all possible corresponding ion configurations. All templates are compared with the column sum data set using the comparison function
@@ Appendix D Mass spectrometry
A radio frequency (RF) “tickling” signal with an amplitude of 30 mV modulates the DC voltage on the endcap electrodes, which defines the axial confinement, and excites the COM motion of the ions which peaks when on resonance with the COM motional frequency. Due to Doppler broadening, the ions fluoresce less when their motional state is excitedWineland, Drullinger, and Walls (1978). The COM motional frequency can then be determined by scanning the RF tickling frequency to minimize the fluorescence.
In crystals with mixed ion species, the COM motional frequency depends not only on the different ion masses but also on the specific ion crystal configurationHome (2013). For instance, the two ion configurations (Ca, Ca, CaOH) and (Ca, CaOH, Ca) have a different COM motional frequency, with a deviation of 1.7 kHz in our system. When applying an RF signal which is close to the COM motional frequency of the current ion configuration, the ions reconfigure their respective positions due to the induced motion. After the reconfiguration, the RF signal is no longer resonant with the COM motion which makes it difficult to perform a precise measurement of the resonance frequency. Thus, we perform measurements with crystals consisting of exactly two trapped ions, as there exists only a single COM motional frequency due to symmetry. However, the dissociation measurements are performed with larger ion crystals to speed up the CaOH generation process. In order to prevent ion loss during the tickling scan, a Doppler cooling cycle of about is applied to the ions with the RF signal being turned off after each tickling measurement in the scan. For every single Ca and molecular ion pair, we performed and summed up to five measurement runs.
@@
The fitted COM motional frequencies from the 65 Ca and molecular ion pairs and the corresponding reference COM motional frequencies of Ca gave masses for the molecular ions ranging from amu. This confirms that all measured 65 molecular ions were CaOH.
In order to estimate the statistical uncertainty of the measured COM motional frequency of trapped atomic Ca ions, we evaluated 34 data sets taken over the course of hours on two different days. For a voltage of 205 V applied to the endcap electrodes, a COM motional frequency of 526.8 kHz with a standard deviation of 200 Hz was determined. Since the errors in the fitted COM motional frequency of all 65 processed datasets of Ca and molecular ion pairs and all reference datasets are all smaller than 200 Hz, this value is utilized to obtain an upper limit for the uncertainty of COM motional frequencies. This results in an estimated uncertainty in the derived molecular ion masses of 0.1 amu.

### v2

@@ Acknowledgements
This research was funded by ERC-2020-STG 948893, ESQ Discovery project SDEF: State-dependent force spectroscopy for trapped ions, FWF 1000 Ideas project TAI-798, and the Foundation for Polish Science within the First Team programme. We acknowledge Poland’s high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational project PLG/2023/016115. The authors would like to acknowledge AQT for support with the ion trap and vacuum pumps, in particular Daniel Nigg and Georg Jacob; Markus Teller for assistance with vacuum work; electronics workshop technicians Kilian Prokop and Wolfgang Kuen for lab infrastructure support; Milan Oncak for discussions on quantum chemistry calculations; Christian Marciniak for general discussions; Marco Valentini for Ca+ lasers implementation; and the Quantum Optics & Spectroscopy Group and associated ion trapping groups at the Universität Innsbruck for general assistance.
Author Contributions Z.W., S.W., B.F., and P.S. designed the experiment. Z.W., S.W., and B.F. carried out the dissociation experiment and analyzed the data. M.I.M., E.M., and B.F. performed the characterization of the OPA system. P.G. and M.T. performed electronic structure calculations. Z.W., S.W., B.F. and M.T. contributed to the manuscript. Z.W., S.W., V.P., M.I.M., E.M., G.M., R.N., B.F., and P.S. contributed to the experimental setup. All authors reviewed the manuscript. B.F. and P.S. supervised the project.
@@
The data that support the findings of this study are openly available at http://doi.org/10.5281/zenodo.11109790.
@@ A.3 Memory of the photodissociation process
We can therefore also estimate the mean dissociation time using the estimated dissociation probability for fixed times. These probabilities are fitted to the exponential distribution CDF as indicated in Fig. 8, e.g., for nm and nm. We estimate the photodissociation cross sections via this analysis for all data and compare them to the cumulative dissociation time, as shown in Fig. 9. We find that the cross sections obtained from both methods agree within the uncertainties which indicates that the single- and two-photon dissociation processes are memoryless.
@@ Appendix B Experimental details
B.1 Ca+ Lasers and trap parameters
@@ Appendix C Ion crystal configuration identification
The photodissociation experiment requires real-time evaluation of the ion crystal configuration such that the number of fluorescing (bright) Ca+ and non-fluorescing (dark) molecular ions can be monitored. Knowledge of the ion crystal configuration allows a script to run automatic measurements and take actions such as (re)loading the trap or emptying it if more than one dark ion is created. The ion crystal configuration is identified using solely image data of the ions taken by an imaging system with magnification using an Andor iXon+ 885 camera with a pixel size of 8 µm. For this process, a region of interest (ROI) in the image centered around the ions is converted into two normalized data sets: the column sum, where the ROI is summed along the columns of the image array; and the row sum, where the summation is done along the rows (see Fig. 12). All analysis to identify the ion crystal configuration is based on these two data sets.
@@
The number of bright ions is counted from the column sum via a peak finder algorithm, which can determine the number and positions of bright ions without requiring a fluorescence-based threshold. This algorithm extracts and sorts local maxima of column sum over subsets of length comparable but smaller than the ion separation. From the largest amplitude difference and their amplitudes, the algorithm classifies peaks into two distinct groups: large-amplitude peaks originating from bright ions and small-amplitude peaks due to noise. Knowledge of the positions of the bright ion peaks can be utilized to identify the presence of dark ions for, e.g., determining the trap center position, as it requires an ion crystal of solely bright ions.
In order to count dark ions, a template fitting procedure was performed to identify the ion crystal configuration. In this process, template sets for different ion numbers and configurations are compared to the column sum data set. Template sets are generated by the following scheme. The desired ion configuration in a template for ions is given by the vector . The value of is assigned to for a bright Ca+ ion and for a dark molecular ion, for , e.g., the ion crystal configuration template which matches the row sum shown in Fig. 12 is . Then the sum of Gaussian peaks
@@
Here denotes the numerically calculated positionsJames (1998) of the ions with respect to the calibrated trap center position , is the waist of the Gaussian peaks (taken from the fit presented in Fig. 12), and is an offset of the trap center position. The start and end positions, and , of the -th binning interval are chosen such that the binning and length of the template match those of the measured column sum data set. is the binned template of ions for the configuration . To find the particular ion crystal configuration, a set of templates is created for all reasonable dark and bright ion numbers and all possible corresponding ion configurations. All templates are compared with the column sum data set using the comparison function
@@ Appendix D Mass spectrometry
To determine the molecular mass, a radio frequency (RF) “tickling” signal with an amplitude of 30 mV modulates the DC voltage on the endcap electrodes which defines the axial confinement. The signal excites the COM motion of the ions which peaks when the signal frequency is on resonance with the COM motional frequency. Due to Doppler broadening, the ions fluoresce less when their motional state is excitedWineland, Drullinger, and Walls (1978). The COM motional frequency can then be determined by scanning the RF tickling frequency to minimize the fluorescence.
In crystals with mixed ion species, the COM motional frequency depends not only on the different ion masses but also on the specific ion crystal configurationHome (2013). For instance, the two ion configurations (Ca+, Ca+, CaOH+) and (Ca+, CaOH+, Ca+) have a different COM motional frequency, with a deviation of 1.7 kHz in our system. When applying an RF signal which is close to the COM motional frequency of the current ion configuration, the ions reconfigure their respective positions due to the induced motion. After the reconfiguration, the RF signal is no longer resonant with the COM motion which makes it difficult to perform a precise measurement of the resonance frequency. Thus, we perform measurements with crystals consisting of exactly two trapped ions, as there exists only a single COM motional frequency due to symmetry. However, the dissociation measurements are performed with larger ion crystals to speed up the CaOH+ generation process. To prevent ion loss during the tickling scan, a Doppler cooling cycle of about is applied to the ions with the RF signal being turned off after each tickling measurement in the scan. For every single Ca+ and molecular ion pair, we performed and summed up to five measurement runs.
@@
The fitted COM motional frequencies from 65 Ca+ and molecular ion pairs and the corresponding reference COM motional frequencies of Ca+ gave masses for the molecular ions ranging from amu. This confirms that all measured 65 molecular ions were CaOH+.
In order to estimate the statistical uncertainty of the measured COM motional frequency of trapped atomic Ca+ ions, we evaluated 34 data sets taken over the course of hours on two different days. For a voltage of 205 V applied to the endcap electrodes, a COM motional frequency of 526.8 kHz with a standard deviation of 200 Hz was determined. Since the errors in the fitted COM motional frequency of all 65 processed datasets of Ca+ and molecular ion pairs and all reference datasets are all smaller than 200 Hz, this value is utilized to obtain an upper limit for the uncertainty of COM motional frequencies. This results in an estimated uncertainty in the derived molecular ion masses of 0.1 amu.